You are going to make a box out of cardboard. The box needs to measure 10 cm by 5 cm by 5 cm; 10% of the final product will be wasted in the cutting process. How much cardboard do you need?
Answer by checkley77(12569)

The final surface area of the box is:
10*5=50 for the base.
2*10*5=100 for the 2 long sides.
2*5*5=50 for the 2 short sides.
50+100+50=200 cm^2 for the total surface.
.10*200=20 cm^2 will be wasted?
Graphing the problem this answer doesn't make sense. By cutting 4 5x5 corners this accounts for 100 cm^2 wasted out of a 15x20=300 cm^2 piece of cardboard.
Here the waste is 100 cm^2 out of 300 cm^2 (33%) original size or 100 cm^2 out of 200 cm^2 (50%) for the box.